# SweetTokenizer: Semantic-Aware Spatial-Temporal Tokenizer for Compact Visual Discretization

### v2

@@ 4.5 Visual Semantic Comprehension
To evaluate the semantic capabilities of SweetTokenizer, we conducted experiments on few-shot image classification and video action recognition tasks. In both experiments, we initially extracted visual tokens using SweetTokenizer and transformed them into natural language words via our LLM codebook. Subsequently, we employed CLIP to compute the similarity between the visual inputs and text embeddings. The top 21 tokens with the highest similarity were selected to form a prompt for prediction using the Qwen LLM. For the image classification task, we adhered to the V2L protocol [59], comparing SweetTokenizer against other language-based visual tokenizers SPAE [54] and V2L. In the video action recognition task, we used ARN [58] and HF-AR [25] as baselines. The results in Table 6 indicate that SweetTokenizer achieved an accuracy of 90.8% on the miniImageNet dataset, surpassing SPAE 89.9% and V2L 87.6%. On the UCF-101 dataset, SweetTokenizer attain an average accuracy of 90.1%, outperforming ARN 83.1% and HF-AR 86.4%. These findings demonstrate SweetTokenizer robust semantic understanding and superior performance in both image and video tasks.
@@ 5 Conclusions
We present SweetTokenizer, an efficient visual tokenization framework that compresses spatial and temporal information through the cross-attention query autoencoder. Combined with language-based latent spaces and progressive training schedule, SweetTokenizer reduces token count by 4$\times$ for video data while maintaining high reconstruction fidelity. Our approach offers a compact representation of visual data, making it well-suited for downstream tasks such as visual generation, visual recognition, marking a significant step in efficient visual tokenization.
@@
$CQAE_{s,t}$ Cross-attention query autoencoder
@@
$\Omega_{s,t,m}$ Continuous latent query tokens
$z_{s,t,m}$ Continuous latent query token
$\hat{\Omega}_{s,t}$ Quantized latent query tokens
$\hat{z}_{s,t}$ Quantized latent query token
@@
$\hat{v}_{s,t}$ Discrete visual feature
$h(\cdot)$ Hidden unit
@@
$\mathcal{P}$ Projector network
@@
$\Psi_{e,d}(\cdot)$ Pretrained encoder & decoder
$\mathcal{L}_{c}$ Cross-Entropy loss
$\mathcal{L}_{vq}$ Commitment loss
$\mathcal{L}_{m}$ Mean-squared loss
$\mathcal{L}_{g}$ Gan loss
$\mathcal{L}_{p}$ Perceptual loss
@@ 6.1 Model Implementation Details
The tokenizer is composed of an encoder $\mathcal{E}$, decoder $\mathcal{D}$, and latent quantizer $\mathcal{Q}$. The tokenizer takes a video clip of 17 consecutive frames with a resolution of 256 $\times$ 256 with the elements normalized to $[-0.5,0.5]$ as input. Then the video clip will be patchified to a resolution of 1 $\times$ 32 $\times$ 32 spatial feature and 4 $\times$ 32 $\times$ 32 temporal feature as illustrated in the main paper. The encoder $\mathcal{E}$ and decoder $\mathcal{D}$ in our tokenizer are both composed of 8 $CQAE_{s}$ modules and 4 $CQAE_{t}$ modules with 512 hidden states and 8 attention heads. Each modules consists of self-attention, feed-forward and cross-attention layers. Before the attention computation, the visual features will be reshaped into $[(BT)\times(HW)\times D]$ and $[(BHW)\times T\times D]$ for $CQAE_{s}$ and $CQAE_{t}$ modules, respectively. The encoder $\mathcal{E}$ generates 256 spatial and 1024 temporal continuous latent tokens. These tokens are then passed to the quantizer $\mathcal{Q}$, which produces the quantized spatial and temporal latent tokens. The quantizer $\mathcal{Q}$ is composed of a spatial and temporal codebook and a GCN with two hidden layers with hidden dimension of 512 as the projector network $\mathcal{P}$. To improve the training stability of the visual tokenizer, we adopt exponential moving average (EMA) updates with weight of 0.999 following [56].
@@
Tokenizer #Params #Tokens FID $\downarrow$
VQVAE-2 [32] 13.5B 1024 31.11
VQGAN [11] 1.4B 1024 74.3
RQ-Transformer [26] 821M 1024 13.11
ViT-VQGAN [46] 650M 1024 8.81
OmniTokenizer [45] 650M 1024 7.45
SweetTokenizer 650M 256 5.48
Table 8: Comparison of class-conditional generation results on ImageNet-1K. Each image is of resolution of 256 $\times$ 256. We only compare autoregressive generators (AR) as baselines. “#Params” refers the parameter number of generator.
@@
Stage 1 Stage 2 Stage 3
LLM init VideoGPT - - -
Optimizer AdamW AdamW AdamW AdamW
@@
Batch size per GPU 4 32 8 12
Peak learning rate $1e^{-4}$ $1e^{-4}$ $1e^{-4}$ $1e^{-4}$
Discriminator peak learning rate - - - $1e^{-4}$
Learning rate schedule Cosine Cosine Cosine Cosine
Training steps 1000K 1000K 500K 500K
Discriminator start steps - - - 20K
Warm-up steps 10K 10K 10K 10K
Weight decay 0.03 $1e^{-4}$ $1e^{-4}$ $1e^{-4}$
Numerical precision float16 float16 float16 bfloat16
Table 9: The detailed training hyperparameters of SweetTokenizer.
@@ 6.4 Training Settings
The detailed training hyper-parameter settings for SweetTokenizer are reported in Table 9.
@@ 7 Additional Results
7.1 Generation results for ImageNet
Following the OmniTokenizer protocol [45], we train a VideoGPT-based generator on ImageNet-1K. As shown in Table 8, SweetTokenizer consistently outperforms all baselines in terms of FID. With fewer tokens, SweetTokenizer achieves 5.48 FID, outperforming OmniTokenizer by 26.4 %. These results align with findings on UCF-101, reinforcing that fewer tokens lead to better generation performance.
@@ 7.2 More Visualizations
Fig 8 and Fig 7 visualize the reconstruction results for the UCF-101 and K-600 datasets. The pixel-level differences between ground truth and model are shown, with brighter areas indicating greater disparity and darker areas reflecting consistency. As shown, SweetTokenizer exhibits fewer reconstruction differences compared to OmniTokenizer, demonstrating its superior performance.
Fig 8 and Fig 9 visualize the reconstruction and generation results of SweetTokenizer on ImageNet-1K. For reconstruction, differences between models are highlighted in red blocks, with details shown in green blocks. Clearly, SweetTokenizer outperforms all baselines by a significant margin.
Finally, we visualize the words from our LLM codebook in Fig 10, based on few-shot video action recognition tasks on the UCF-101 dataset. We use adjectives, nouns, adverbs, and verbs as prompts to Qwen LLM for action prediction. Green and orange indicate meaningful words, while red marks meaningless ones. The visualization shows that correct verb words consistently lead to accurate predictions, even when other words are irrelevant, highlighting the importance of our adverb and verb LLM codebook modules for video action recognition.
Figure 6: Comparison of the reconstruction results of OmniTokenizer and SweetTokenizer on UCF-101 dataset, where ”Diff” represents the pixel difference between the ground truth and the models.
Figure 7: Comparison of the reconstruction results of OmniTokenizer and SweetTokenizer on K-600 dataset, where ”Diff” represents the pixel difference between the ground truth and the models.
Figure 8: Comparison of the reconstruction results of TiTok, OmniTokenizer, and SweetTokenizer on ImageNet-1K dataset. Differences are selected by the red blocks and highlighted in the grean blocks.
Figure 9: Class-conditional image generation results using SweetTokenizer, on ImageNet-1k dataset.
@@ 8 Limitations
Our tokenizer is not suitable for tasks requiring precise semantic understanding, like VQA, because the LLM codebook is trained in an unsupervised manner. Without additional constraints, such as contrastive learning between image features from Qwen-VLM and text embeddings in our codebook, aligning the image and text domains is challenging. A promising direction for future work is to enhance SweetTokenizer into a semantically strong tokenizer by contrastive learning.

### v3

@@ 4.5 Visual Semantic Comprehension
To evaluate the semantic capabilities of SweetTok, we conducted experiments on few-shot image classification and video action recognition tasks. In both experiments, we initially extracted visual tokens using SweetTok and transformed them into natural language words via our language-based codebook. Subsequently, we employed CLIP to compute the similarity between the visual inputs and text embeddings. The top 21 tokens with the highest similarity were selected to form a prompt for prediction using the Qwen LLM. For the image classification task, we adhered to the V2L protocol [54], comparing SweetTok against other language-based visual tokenizers SPAE [49] and V2L. In the video action recognition task, we used ARN [53] and HF-AR [20] as baselines. The results in Table 6 indicate that SweetTok achieved an accuracy of 90.8% on the miniImageNet dataset, surpassing SPAE 89.9% and V2L 87.6%. On the UCF-101 dataset, SweetTok attain an average accuracy of 90.1%, outperforming ARN 83.1% and HF-AR 86.4%. These findings demonstrate SweetTok robust semantic understanding and superior performance in both image and video tasks. We also visualize the effectiveness of our MLC spatial-temporal words for video understanding in the supplementary.
Methods rFVD $\downarrow$
Baseline (w/o LC) 29.45
$+$ LC 24.80
$+$ MLC (SweetTok) 20.46
$+$ CLIP [27]-based MLC (SweetTok) 20.46
$+$ Qwen [1]-based MLC 20.12
Table 5: Ablation study of different codebooks, including vanilla language codebook (LC), motion-enhanced language codebook (MLC) and more advanced pre-trained language codebook.
@@ 5 Conclusions
We present SweetTok, an efficient video tokenization framework that compresses spatial and temporal information through the decoupled query autoencoder. Combined with motion-enhanced language codebook, SweetTok reduces token count for video data more effectively, achieving higher reconstruction fidelity compared to previous state-of-the-arts. Our approach offers a compact representation of video data, making it well-suited for downstream tasks such as video generation, understanding, marking a significant step in efficient video tokenization.
@@
$DQAE_{s,t}$ Decoupled query autoencoder
$\mathcal{P}$ Patchify Module
@@
$\mathbf{Q}_{s,t,m}$ Continuous latent query tokens
$\mathbf{Z}_{\mathbf{Q}_{s,t,m}}$ Continuous latent query token
$\tilde{\mathbf{Q}}_{s,t}$ Quantized latent query tokens
$\tilde{\mathbf{Z}}_{\mathbf{Q}_{s,t,m}}$ Quantized latent query token
@@
$\tilde{v}_{s,t}$ Discrete visual feature
@@
$\mathcal{F}$ Projector network
@@ 6.1 Model Implementation Details
The tokenizer is composed of an encoder $\mathcal{E}$, decoder $\mathcal{D}$, and latent quantizer $\mathcal{Q}$. The tokenizer takes a video clip of 17 consecutive frames with a resolution of 256 $\times$ 256 with the elements normalized to $[-0.5,0.5]$ as input. Then the video clip will be patchified to a resolution of 1 $\times$ 32 $\times$ 32 spatial feature and 4 $\times$ 32 $\times$ 32 temporal feature as illustrated in the main paper. The encoder $\mathcal{E}$ and decoder $\mathcal{D}$ in our tokenizer are both composed of 8 $DQAE_{s}$ modules and 4 $DQAE_{t}$ modules with 512 hidden states and 8 attention heads. Each modules consists of self-attention, feed-forward and cross-attention layers. Before the attention computation, the visual features will be reshaped into $[(BT)\times(HW)\times D]$ and $[(BHW)\times T\times D]$ for $DQAE_{s}$ and $DQAE_{t}$ modules, respectively. The encoder $\mathcal{E}$ generates 256 spatial and 1024 temporal continuous latent tokens. These tokens are then passed to the quantizer $\mathcal{Q}$, which produces the quantized spatial and temporal latent tokens. The quantizer $\mathcal{Q}$ is composed of a spatial and temporal codebook and a GCN with two hidden layers with hidden dimension of 512 as the projector network $\mathcal{F}$. To improve the training stability of the visual tokenizer, we adopt exponential moving average (EMA) updates with weight of 0.999 following [51].
@@
Tokenizer PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
UCF-101
LARP [39] 27.88 - 0.085
SweetTok 29.27 0.7766 0.070
ImageNet
LlamaGen-16 [31] 20.79 0.675 -
TokenFlow [25] 21.41 0.687 -
LlamaGen-8 [31] 24.45 0.813 -
SweetTok 30.23 0.826 0.068
Table 8: More evaluation results on UCF-101 and ImageNet.
@@
Image Finetune Video Training
LLM init VideoGPT - -
Optimizer AdamW AdamW AdamW
@@
Batch size per GPU 4 32 12
Peak learning rate $1e^{-4}$ $1e^{-4}$ $1e^{-4}$
Discriminator peak learning rate - - $1e^{-4}$
Learning rate schedule Cosine Cosine Cosine
Training steps 1000K 500K 1000K
Discriminator start steps - - 20K
Warm-up steps 10K 10K 10K
Weight decay 0.03 $1e^{-4}$ $1e^{-4}$
Numerical precision float16 float16 bfloat16
Table 9: The detailed training hyperparameters of SweetTok.
@@ 6.4 Training Settings
The detailed training hyper-parameter settings for SweetTok are reported in Table 9. During video training, we train the first 500K steps using proxy code following [51] to accelerate training. The remaining 500K is trained without proxy code.
@@ 7 Additional Results
7.1 More evaluation metrics
We assess SweetTok using additional metrics: PSNR, SSIM, and LPIPS. As shown in Table 8, SweetTok outperforms all baselines on both video and image datasets, further validating the superiority of our model design.
@@ 7.2 More Visualizations
Fig 9 and Fig 8 visualize the reconstruction results for the UCF-101 and K-600 datasets. The pixel-level differences between ground truth and model are shown, with brighter areas indicating greater disparity and darker areas reflecting consistency. As shown, SweetTok exhibits fewer reconstruction differences compared to OmniTokenizer, demonstrating its superior performance.
Fig 9 visualize the reconstruction results of SweetTok on ImageNet-1K. For reconstruction, differences between models are highlighted in red blocks, with details shown in green blocks. Clearly, SweetTok outperforms all baselines by a significant margin.
Finally, we visualize the words from our MLC in Fig 10, based on few-shot video action recognition tasks on the UCF-101 dataset. We use adjectives, nouns, adverbs, and verbs as prompts to Qwen LLM for action prediction. Green and orange indicate meaningful words, while red marks meaningless ones. The visualization shows that correct verb words consistently lead to accurate predictions, even when other words are irrelevant, highlighting the importance of our MLC modules for video action recognition.
Figure 7: Comparison of the reconstruction results of OmniTokenizer and SweetTok on UCF-101 dataset, where ”Diff” represents the pixel difference between the ground truth and the models.
Figure 8: Comparison of the reconstruction results of OmniTokenizer and SweetTok on K-600 dataset, where ”Diff” represents the pixel difference between the ground truth and the models.
Figure 9: Comparison of the reconstruction results of TiTok, OmniTokenizer, and SweetTok on ImageNet-1K dataset. Differences are selected by the red blocks and highlighted in the grean blocks.
@@ 8 Limitations
Our tokenizer is not suitable for tasks requiring precise semantic understanding, like VQA, because the MLC is trained in an unsupervised manner. Without additional constraints, such as contrastive learning between image features from Qwen-VLM and text embeddings in our codebook, aligning the image and text domains is challenging. A promising direction for future work is to enhance SweetTok into a semantically strong tokenizer by contrastive learning.